If the mean of random variable X is 2.5, then the mean of 2X−3 is

a)5 b)2 c)7 d)none of the above
Answer #1

Mean of X, E(X) = 2.5(X+k) =

Then, mean of 2X-3 = E(2X-3) = 2*E(X) - 3, since E(cX) = c*E(X) if c is a constant and E(X+k) = E(X)+k if k is a constant

= (2*2.5) - 3 = 2 (Option b)